OpenStudy (unklerhaukus):

Simplify\[\sqrt{1000}\]

OpenStudy (anonymous):

31.622

OpenStudy (unklerhaukus):

but how

OpenStudy (anonymous):

Its kinda hard to do without of calculator, can't really explain sorry :\

OpenStudy (sohailiftikhar):

umm ok divide this into different numbers like \[\sqrt{500 \times 2 }=\sqrt{250 \times 4}=\sqrt{125 \times 2 \times 4}=\sqrt{5^3 \times 2 \times }\]

OpenStudy (sohailiftikhar):

\[(5)(2)\sqrt{5 \times 2}\]

OpenStudy (anonymous):

God damn muslim tryna overdo me xD

OpenStudy (sohailiftikhar):

\[10\sqrt{5}.\sqrt{2}=(10)(1.4)\sqrt{5}\] because sqrt2=1.4

OpenStudy (unklerhaukus):

hmm 10.√5.√2, that seems best

OpenStudy (sohailiftikhar):

now it is simple enough \[14\sqrt{5}\]

OpenStudy (sohailiftikhar):

yeah! :)

OpenStudy (anonymous):

Oh, sht, see I gave you the wrong answer, cause I thought u just meant find the sqrt of 1000 xD, dont blame me im new to this thing

OpenStudy (sohailiftikhar):

so ? close this question dude :)